OpenStudy (sonyalee77):

If given f(x) = x + 4 and g(x) = –2x^2 + 3x – 1, what is (f – g)(x)? A. (f – g)(x) = –2x^2 + 4x – 3 B. (f – g)(x) = –2x^2 – 2x + 5 C. (f – g)(x) = 2x^2 + 2x + 5 D. (f – g)(x) = 2x^2 – 2x + 5

OpenStudy (sonyalee77):

@dumbcow @Euler271

OpenStudy (dumbcow):

f - g = (x+4) - (-2x^2+3x-1) = 2x^2 -2x +5

OpenStudy (sonyalee77):

thanx i'm on a timer could u help me with a few more?

OpenStudy (dumbcow):

haha why would you say that....thats a dead give away that you are taking some kind of online test

OpenStudy (sonyalee77):

yes i am its just tht some of this stuff i dnt remember learning in my lesson. I take E2020